OpenStudy (anonymous):

what is the range and domain of y = log base of 5 x

OpenStudy (anonymous):

x can't be negative or 0. So the domain is (0, infinity).

OpenStudy (anonymous):

& the range is (-infinity, +infinity) because, no matter which real number you substitute for y, you'll get a corresponding x such that \(\Large5^y=x\)

OpenStudy (anonymous):

@sahura do you want me to explain it furthur?

OpenStudy (anonymous):

thank you so much , u dont have to explain it furthur bt i also need help with this other problem : how do i find the inverse of y= log [base 4] x

OpenStudy (anonymous):

this statement will be your friend for all questions like this: The logarithm of a number w.r.t a base is the index to which the base must be raised in order to get the number. This statement should allow you to write x in terms of y.

OpenStudy (anonymous):

@rajathsbhat, im sorry, i still dont understand how to solve it

OpenStudy (anonymous):

could u please explain further how to solve what is the inverse of y = log base of 4 x

OpenStudy (anonymous):

so the link is saying that 4x is the inverse , so whats the answer?

OpenStudy (anonymous):

you see, by definition (which i have mentioned in the previous reply), \[\Large y=\log _4x\Rightarrow x=4^{y}\] This implies that 'x' is the inverse of 'y'. And what is that 'x'? It is \(4^{y}\)

OpenStudy (anonymous):

Put simply, \(x=4^{y}\) is the inverse function of \(y=\log _4x\)

OpenStudy (anonymous):

oh now i understand, thank you so much for your help :)

OpenStudy (anonymous):

:)